# Real-Time Recurrent Reinforcement Learning

### v2

@@ Discussion
We introduced real-time recurrent reinforcement learning (RTRRL), a new approach to solving discrete and continuous control tasks for POMDPs, in a biologically plausible fashion. RTRRL consists of three parts: 1) A Meta-RL RNN architecture, implementing on its own a nested RL algorithm, 2) A TD($\lambda$) actor-critic algorithm exploiting temporal difference and dutch eligibility traces, to train the weights of the Meta-RL network, and 3) A biologically plausible, random-feedback local-online (RFLO) real-time optimization algorithm, for computing the weight gradients.
We compared RTRRL with popular, but biologically implausible RL algorithms that use BPTT or RTRL for gradient-computation. Our results show, that using approximate gradients as in RFLO, still finds satisfactory solutions, and in some cases, it even improves on the state of the art. Particularly, RTRRL outperformed PPO with BPTT, when solving tasks requiring exploration in unfavourable environments.
Although the results presented in this paper are empirically convincing, some limitations have to be discussed. The algorithm generally suffers from a larger variance due to an effective batch size of 1. A possible improvement of RTRRL might be achieved, by using batched experience replay, which has become a standard in RL. However, this has to be done in a way, that preserves biological plausibility.
RTRRL is grounded in neuroscience and adequately explains how biological neural networks learn to act in unknown environments. The network structure resembles the interplay of dorsal and ventral striatum of the basal ganglia, featuring global RPEs found in dopaminergic pathways projecting from the ventral tegmental area and the substantia nigra zona compacta to the striatum and cortex (Wang et al. 2018). The role of dopamine as RPE was established experimentally by Wise (2004) who showed that dopamine is release upon receiving an unexpected reward, reinforcing the recent behavior. However, dopamine is also released in response to a conditioned stimulus, before receiving the actual reward. If then the expected reward is absent, dopamine levels will drop below baseline - a negative reinforcement signal. Dopaminergic synapses are usually located at the dendritic stems of glutamate synapses (Kandel 2013) and can therefore effectively mediate synaptic plasticity.
More specifically, the ventral striatum would correspond to the critic in RTRRL and the dorsal striatum to the actor, with dopamine axons targeting both the ventral and dorsal compartmens (Sutton and Barto 2018). The axonal tree of dopaminergic synapses is represented by the backward weights in RTRRL. Dopamine subsequently encodes the TD-error as RPE, which is used to update both the actor and the critic. RTRRL can therefore be seen as a model of reward-based learning taking place in the human brain.
Appendix A Derivation of update equations
@@
Appendix B Implementation Details
@@
Actor learning rate $\alpha_{A}$ 1e-2
Critic learning rate $\alpha_{C}$ 1.0
RNN learning rate $\alpha_{R}$ 1e-3
@@
Table 2: Hyperparameters of RTRRL and values used.
Algorithm 2 Real-Time Recurrent Reinforcement Learning with accumulate eligibility traces and RTRL
@@
0:  Linear value function: $\hat{v}(h,\theta_{C})$
@@
  $\theta_{A},\theta_{C},\theta_{R}\leftarrow$ Randomly initialize parameters
  if RFLO then
     $B_{A},B_{C}\leftarrow$ Randomly initialize feedback matrices
  end if
  $h,e_{A},e_{C},e_{R}\leftarrow\mathbf{0}$
  $o\leftarrow$ Reset Environment
  $h,\hat{J}\leftarrow$RNN$([o,\mathbf{0},0],h,\theta_{R})$
  $v\leftarrow\hat{v}(h,\theta_{C})$
  while not done do
     logits $\leftarrow\pi(h,\theta_{A})$
     $a\leftarrow$ Sample(logits)
     $o,R\leftarrow$ Take action $a$ in Environment
     $h^{\prime},\hat{J}^{\prime}\leftarrow$RNN($[o,a,r],h,\hat{J},\theta_{R}$)
     $v^{\prime}\leftarrow\hat{v}(h^{\prime},\theta_{C})$
     $\delta\leftarrow R+\gamma v^{\prime}-v$
     $e_{C}\leftarrow\gamma\lambda_{C}e_{C}+\nabla_{\theta_{C}}\hat{v}$
[…]
     if RTRL then
[…]
     else if RFLO then
        $g_{C}\leftarrow B_{C}\mathbf{1}$
[…]
        $e_{R}\leftarrow\gamma\lambda_{R}e_{R}+g_{C}+g_{A}$
     end if
     $\theta_{C}\leftarrow\theta_{C}+\alpha_{C}\delta e_{C}$
     $\theta_{A}\leftarrow\theta_{A}+\alpha_{A}\delta e_{A}$
     $\theta_{R}\leftarrow\theta_{R}+\alpha_{R}\hat{J}\delta e_{R}$
     $v\leftarrow v^{\prime},\quad h\leftarrow h^{\prime},\quad\hat{J}\leftarrow\hat{J}^{\prime}$
  end while
Neuron Model.
Figure 7: CT-RNNs are solving an Ordinary Differential Equation. In general, any solver may be used. When using forward Euler, $dt=k^{-1}$ is a hyperparameter that determines the number of solver steps and subsequently the accuracy of the solution.
The simplified CT-RNN outlined in the paper is taken from (Murray 2019). Our code allows for increasing the number of steps $k=dt^{-1}$ when solving the underlying Ordinary Differential Equation with the forward Euler method. More steps lead to a more expressive model meaning you can get away with fewer neurons, but also to increased computational complexity. In our experiments we kept $k=1$ for simplicity.
Figure 8: Evaluation rewards for environments from the brax package that where masked as described in section Experiments to make them POMDPs. Shown are the best run for RTRRL and the tuned baseline shipped with the package.
Appendix C More experiments
Memory Length
We tested the memory capacity of our RTRRL models by learning to remember the state of a bit for an extended number of steps. The experiment is explained in section Experiments. Figure 4 shows the rewards of 10 runs each as boxplots. We also included a naive approximation to RTRL we call LocalMSE that only uses the immediate Jacobian for updates. Looking at the boxplot we can clearly see that RTRL outperforms the other methods and that RFLO is situated at a middle ground between LocalMSE and RTRL in terms of memory capacity.
Physics Simulation
For this set of experiments we created POMDPs with continuous actions by masking the observations of three different brax environments. For each environment, we ran hyperparameter tuning for at least 10 hours and picked the best performing run. Figure 8 shows the evaluation rewards achieved throughout the learning process for RTRRL and the PPO baseline that was provided by the package.
Appendix D Biological implausibility of BPTT
Here, we elaborate further on the objections to biological plausibility of Backpropagation-Through Time.
@@ Appendix D Biological implausibility of BPTT
Appendix E Biological Interpretation of RTRRL
Real-Time Recurrent Reinforcement Learning is a model of goal-directed behaviour learning and training of reflexes in animals, located in the human basal ganglia (Rusu and Pennartz ). It comprises a scalar-valued global reward signal that can be interpreted as dopaminergic synapses, a value estimator (critic) and motor output (actor).
@@ Appendix E Biological Interpretation of RTRRL
There is ample evidence that certain brain structures are implementing actor-critic methods. The striatum is involved in motor and action planning, decision-making, motivation, reinforcement and reward perception. It is also heavily innervated by dopamine axons coming from the VTA and the SNc. It is speculated that dopamine release in the ventral striatum ”energizes the next response” while it acts by stamping in the procedural memory trace in the dorsal striatum ”establishing and maintaining procedural habit structures” (Wise 2004). Subsequently, the ventral striatum would correspond to the critic and the dorsal striatum to the actor of an RL algorithm (Sutton and Barto 2018). Dopamine would then correspond to the TD-error which is used to update both the actor and the critic. As dopamine neuron axons target both the ventral and dorsal striatum, and dopamine appears to be critical for synaptic plasticity, the similarities are evident. Furthermore, the TD-error and dopamine levels are both encoding the RPE: they are high whenever an unexpected reward is received and they are low (or negative in case of the TD-error) when an expected reward does not occur. (Sutton and Barto 2018) These similarities could be beneficial for RL as well as for Neuroscience as advances in either field could lead to new insights that are beneficial to the other.
Appendix F Timing comparison
In order to compare wall-clock time, we did a quick performance test for RTRRL with CT-RNN and PPO with LSTM. In both cases we trained for 5000 steps, with 32 units and a batch size of 1, on a machine with a single GeForce RTX 2070 GPU. For both cases, we repeated the test 3 times and calculated the average time per step of the algorithm. Our results were 7,58 ms / step for PPO-LSTM and 7,49 ms / step for RTRRL-CTRNN. Please note that these results may not carry over to larger model or batch sizes.

### v3

@@ Discussion
We introduced real-time recurrent reinforcement learning (RTRRL), a novel approach to solving discrete and continuous control tasks for POMDPs, in a biologically plausible fashion. We compared RTRRL with PPO, which uses biologically implausible BPTT for gradient-computation. Our results show, that RTRRL with LRUs outperforms PPO consistently when using the same number of neurons. We further found, that using approximate gradients as in RFLO and FA, can still find satisfactory solutions in many cases.
Although the results presented in this paper are empirically convincing, some limitations have to be discussed. The algorithm naturally suffers from a large variance when using a batch size of 1. It moreover needs careful hyperparameter tuning, especially when dealing with continuous actions.
RTRRL is grounded in neuroscience and can adequately explain how biological neural networks learn to act in unknown environments. The network structure resembles the interplay of dorsal and ventral striatum of the basal ganglia, featuring global RPEs found in dopaminergic pathways projecting from the ventral tegmental area and the substantia nigra zona compacta to the striatum and cortex (Wang et al. 2018). The role of dopamine as RPE was established experimentally by Wise (2004) who showed that dopamine is released upon receiving an unexpected reward, reinforcing the recent behavior. If an expected reward is absent, dopamine levels drop below baseline - a negative reinforcement signal. Dopaminergic synapses are usually located at the dendritic stems of glutamate synapses (Kandel 2013) and can therefore effectively mediate synaptic plasticity. More specifically, the ventral striatum would correspond to the critic in RTRRL and the dorsal striatum to the actor, with dopamine axons targeting both the ventral and dorsal compartmens (Sutton and Barto 2018). The axonal tree of dopaminergic synapses is represented by the backward weights in RTRRL. Dopamine subsequently corresponds to the TD-error as RPE, which is used to update both the actor and the critic. RTRRL can therefore be seen as a model of reward-based learning taking place in the human brain.
Finally, an important reason for investigating online training algorithms of neural networks is the promise of energy-efficient neuromorphic hardware. (Zenke and Neftci 2021) The aim is to create integrated circuits that mimic biological neurons. Importantly, neuromorphic algorithms require biologically plausible update rules to enable efficient training.
Acknowledgments
Julian Lemmel is partially supported by the Doctoral College Resilient Embedded Systems (DC-RES) of TU Vienna and the Austrian Research Promotion Agency (FFG) Project grant No. FO999899799. Computational results have been achieved in part using the Vienna Scientific Cluster (VSC).
Appendix A Real-Time Recurrent Reinforcement Learning Appendix
Appendix B Derivation of update equations
@@
Appendix C Implementation Details
Figure 7: Left: RTRRL can be devided into 4 parts that are repeated throughout training. Right: CT-RNNs are solving an Ordinary Differential Equation. In general, any solver may be used. When using forward Euler, $dt=k^{-1}$ is a hyperparameter that determines the number of solver steps and subsequently the accuracy of the solution.
@@ Appendix C Implementation Details
Actor learning rate $\alpha_{A}$ 1e-4
Critic learning rate $\alpha_{C}$ 1e-4
RNN learning rate $\alpha_{R}$ 1e-4
actor RNN trace scale $\eta_{A}$ 1.0
@@
Table 1: Hyperparameters of RTRRL
Neuron Model.
The simplified CT-RNN outlined in the paper is taken from Murray (2019). Our code allows for increasing the number of steps $k=dt^{-1}$ when solving the underlying Ordinary Differential Equation with the forward Euler method. More steps lead to a more expressive model meaning you can get away with fewer neurons, but also to increased computational complexity. In our experiments we kept $k=1$ for simplicity.
Algorithm 2 Real-Time Recurrent Reinforcement Learning with precise backpropagation and entropy regularization
@@
0:  Linear value-function: $\hat{v}(h,\theta_{C})$
@@
1:  $\theta_{A},\theta_{C},\theta_{R}\leftarrow$ Randomly initialize parameters
2:  if RFLO then
3:     $B_{A},B_{C}\leftarrow$ Randomly initialize feedback matrices
4:  end if
5:  $h,e_{A},e_{C},e_{R}\leftarrow\mathbf{0}$
6:  $o\leftarrow$ Reset Environment
7:  $h,\hat{J}\leftarrow$RNN$([o,\mathbf{0},0],h,\theta_{R})$
8:  $v\leftarrow\hat{v}(h,\theta_{C})$
9:  while not done do
10:     logits $\leftarrow\pi(h,\theta_{A})$
11:     $a\leftarrow$ Sample(logits)
12:     $o,R\leftarrow$ Take action $a$ in Environment
13:     $h^{\prime},\hat{J}^{\prime}\leftarrow$RNN($[o,a,r],h,\hat{J},\theta_{R}$)
14:     $v^{\prime}\leftarrow\hat{v}(h^{\prime},\theta_{C})$
15:     $\delta\leftarrow R+\gamma v^{\prime}-v$
16:     $e_{C}\leftarrow\gamma\lambda_{C}e_{C}+\nabla_{\theta_{C}}\hat{v}$
17:     $e_{A}\leftarrow\gamma\lambda_{A}e_{A}+\nabla_{\theta_{A}}\ln\pi[a]$
18:     if RTRL then
19:        $e_{R}\leftarrow\gamma\lambda_{R}e_{R}+\nabla_{\theta_{R}}[\hat{v}+\eta_{\pi}\ln\pi[a]]$
20:        $g_{H,R}\leftarrow\eta_{H}\nabla_{\theta_{R}}H(\pi)$
21:     else if RFLO then
22:        $g_{C}\leftarrow B_{C}\mathbf{1}$
23:        $g_{A}\leftarrow B_{A}\nabla_{h}\ln\pi[a]$
24:        $e_{R}\leftarrow\gamma\lambda_{R}e_{R}+\hat{J}(g_{C}+\eta_{A}g_{A})$
25:        $g_{H,R}\leftarrow\eta_{H}\hat{J}\nabla_{h}H(\pi)$
26:     end if
27:     $\theta_{C}\leftarrow\theta_{C}+\alpha_{C}\delta e_{C}$
28:     $\theta_{A}\leftarrow\theta_{A}+\alpha_{A}(\delta e_{A}+\eta_{H}\nabla_{\theta_{A}}H(\pi))$
29:     $\theta_{R}\leftarrow\theta_{R}+\alpha_{R}(\delta e_{R}+g_{H,R})$
30:     $v\leftarrow v^{\prime},\quad h\leftarrow h^{\prime},\quad\hat{J}\leftarrow\hat{J}^{\prime}$
31:  end while
Appendix D More Experiments
Model Linear TD($\lambda$) RTRL CTRNN RFLO CTRNN PPO CTRNN
Environment
CartPole-vel 454.55$\pm$34.47 500.00$\pm$0.00 500.00$\pm$0.00 112.11$\pm$40.38
CartPole-pos 57.47$\pm$1.38 178.57$\pm$93.04 500.00$\pm$163.17 70.76$\pm$11.42
MetaMaze 28.00$\pm$43.33 34.00$\pm$3.28 40.00$\pm$7.30 10.80$\pm$1.34
DiscountingChain 1.10$\pm$0.04 1.05$\pm$0.04 1.10$\pm$0.05 1.10$\pm$0.00
UmbrellaChain 0.79$\pm$0.05 0.99$\pm$0.15 1.29$\pm$0.07 1.04$\pm$0.12
BernoulliBandit 404.40$\pm$20.51 373.80$\pm$18.20 371.70$\pm$5.17 89.14$\pm$18.54
GaussianBandit 0.01$\pm$5.83 -126.17$\pm$70.60 24.05$\pm$21.43 0.00$\pm$0.00
Reacher 106.66$\pm$6.04 114.45$\pm$9.40 111.06$\pm$26.76 16.42$\pm$0.95
Swimmer 91.26$\pm$7.77 97.57$\pm$4.80 442.54$\pm$225.73 31.06$\pm$2.35
MountCarCont -30.00$\pm$8.94 -40.00$\pm$5.48 -0.03$\pm$0.67 -1104.72$\pm$446.12
MountCarCont-vel -41.88$\pm$16.54 -1.31$\pm$16.14 -20.00$\pm$40.80 -1079.70$\pm$204.52
MountCarCont-pos -46.91$\pm$15.50 0.04$\pm$8.43 -59.54$\pm$43.45 -1071.67$\pm$40.27
MemoryChain-4 0.07$\pm$0.01 1.00$\pm$0.11 1.00$\pm$0.00 1.00$\pm$0.00
MemoryChain-8 0.07$\pm$0.02 0.90$\pm$0.25 1.00$\pm$0.00 1.00$\pm$0.00
MemoryChain-16 0.12$\pm$0.04 0.61$\pm$0.27 1.00$\pm$0.00 1.00$\pm$0.00
MemoryChain-32 0.17$\pm$0.03 0.56$\pm$0.13 0.77$\pm$0.18 1.00$\pm$0.00
MemoryChain-64 0.26$\pm$0.07 0.61$\pm$0.13 0.32$\pm$0.08 1.00$\pm$0.47
DeepSea-4 0.99$\pm$0.00 0.99$\pm$0.44 0.99$\pm$0.00 0.99$\pm$0.00
DeepSea-8 0.99$\pm$0.54 0.99$\pm$0.44 0.99$\pm$0.54 0.99$\pm$0.00
DeepSea-16 -0.00$\pm$0.54 0.99$\pm$0.54 0.99$\pm$0.44 0.00$\pm$0.00
Table 2: Summary of RTRRL experiments on Gymnax for networks with 32 neurons. Numbers reported are the median and standard deviation of the best validation reward achieved throughout training for 5 runs (larger is better). ”RTRRL RFLO” denotes our biologically plausible version of RTRRL.
Model Linear TD($\lambda$) RTRRL RFLO RTRRL LRU PPO CTRNN
Environment
MetaMaze 55.93$\pm$43.33 61.60$\pm$9.24 15.80$\pm$5.02 38.79$\pm$27.03
DiscountingChain 1.09$\pm$0.02 1.10$\pm$0.00 1.10$\pm$0.00 1.32$\pm$0.00
UmbrellaChain 1.19$\pm$0.05 1.72$\pm$0.13 1.20$\pm$0.04 1.04$\pm$0.09
Catch 0.27$\pm$0.08 1.00$\pm$0.00 0.39$\pm$0.83 1.00$\pm$0.00
BernoulliBandit 393.38$\pm$20.51 893.96$\pm$11.35 807.96$\pm$100.19 883.44$\pm$18.72
GaussianBandit nan±$\pm$±nan 61.10$\pm$32.98 37.78$\pm$67.01 0.00$\pm$0.00
PointRobot 1.34$\pm$0.21 5.30$\pm$4.30 0.98$\pm$0.49 1.04$\pm$0.57
Reacher 120.79$\pm$8.69 216.28$\pm$16.90 199.77$\pm$9.26 194.98$\pm$4.84
Swimmer 95.87$\pm$5.94 465.08$\pm$17.47 102.17$\pm$70.34 51.78$\pm$7.13
MemoryChain 0.08$\pm$0.01 1.00$\pm$0.00 nan±$\pm$±nan 1.00$\pm$0.00
DeepSea 0.99$\pm$0.00 0.99$\pm$0.00 0.99$\pm$0.00 0.99$\pm$0.00
StatelessCartPole 0.72$\pm$0.02 1.00$\pm$0.00 1.00$\pm$0.00 0.99$\pm$0.02
NoisyStatelessCartPole 0.50$\pm$0.01 0.99$\pm$0.01 0.93$\pm$0.04 0.59$\pm$0.03
Minesweeper 0.14$\pm$0.02 0.02$\pm$0.05 0.12$\pm$0.01 -0.14$\pm$0.05
MultiArmedBandit 0.25$\pm$0.05 0.39$\pm$0.03 0.38$\pm$0.08 0.19$\pm$0.10
RepeatFirst 0.15$\pm$0.12 -0.22$\pm$0.13 0.05$\pm$0.33 0.03$\pm$0.11
RepeatPrevious -0.35$\pm$0.01 0.54$\pm$0.23 1.00$\pm$0.00 0.93$\pm$0.01
Table 3: Summary of RTRRL experiments including LRU experiments for networks with 32 neurons. Numbers reported are the median and standard deviation of the best validation reward achieved throughout training for 5 runs (larger is better). ”RTRRL RFLO” denotes our biologically plausible version of RTRRL.
In Table 3 we summarize all our results for the experiments explained in the remainder of this section. The column ”RFLO CT-RNN” corresponds to biologically plausible RTRRL as defined above, but we have also included results for RTRRL where RTRL is used in place of RFLO. The values presented in the table are the median and standard-deviation of the best evaluation results, throughout the training of 5 runs per environment and algorithm. Here, evaluation refers to running the environment for 10000 steps, which for example corresponds to 10 episodes for environments with a task duration of 1000 steps. As one can see, RTRRL performs best on average, followed by RTRRL with RTRL, as the second best algorithm.
In the remainder of this section we discuss some additional experiments we conducted that did not make it into the main manuscript of the paper.
Deep Exploration.
Exploration versus exploitation is a trade-off, central to all agents learning a task online. The DeepSea environment included in bsuite (Osband et al. 2020) is tailored to benchmark the ability of RL algorithms to explore in unfavourable environments. The agent is required to explore to reach the goal position albeit receiving negative rewards when moving towards it. We give results for exponentially increasing task length.
Reservoir Computing.
The term reservoir computing refers to using a fixed RNN with trainable linear readout weights, sparing the hassle of exploding or vanishing gradients. Yamashita and Hamagami (2022) employed this technique for solving POMDPs. We conducted experiments to test whether optimizing the RNN in RTRRL is actually beneficial. By keeping the RNN fixed after initialization, training the linear actor and critic only, we create our own echo state network. Figure 9 right shows results for the MemoryChain environment. We conclude that training the RNN does improve performance.
Figure 9: Left: While in most cases it does not make a difference, not using the Meta-RL architecture hampers performance in some cases. Right: Keeping the RNN frozen as in echo state networks leads to significantly worse results. Bottom: Comparing RTRRL with and without biologically plausible propagation of gradients using FA. Some environments such as GaussianBandit and MetaMaze, seem to benefit from it, while others do better with regular backpropagation.
Appendix E Biological implausibility of BPTT
The three main objections for the plausibility of BPTT in biological neural networks are: First, the reliance on shared weights between the forward and the backward synapses; Second, reciprocal error-transport, which means propagating back the errors, without interfering with the neural activity (Bartunov et al. 2018); Third, the concern that BPTT requires storing long sequences of the exact activation for each cell (Lillicrap and Santoro 2019). We rephrase the premises of Bartunov et al. (2018) and Lillicrap and Santoro (2019) as follows. A biologically plausible learning algorithm has to be: (1) Local, up to some low-dimensional reward signal, (2) Online, computing parameter updates and the network outputs in parallel, and (3) Without weight transport, in its error computation, meaning that synapses propagating back error signals must not mirror the strength of forward synapses. We further elaborate on each of these requirements for obtaining biological plausibility in the following.
@@ Appendix E Biological implausibility of BPTT
Appendix F Biological Interpretation of RTRRL
Real-Time Recurrent Reinforcement Learning is a model of goal-directed behaviour learning and training of reflexes in animals, located in the human basal ganglia (Rusu and Pennartz 2020). It comprises a scalar-valued global reward signal that can be interpreted as dopaminergic synapses, a value estimator (critic) and motor output (actor).
@@ Appendix F Biological Interpretation of RTRRL
There is ample evidence that certain brain structures are implementing actor-critic methods. The striatum is involved in motor and action planning, decision-making, motivation, reinforcement and reward perception. It is also heavily innervated by dopamine axons coming from the VTA and the SNc. It is speculated that dopamine release in the ventral striatum ”energizes the next response” while it acts by stamping in the procedural memory trace in the dorsal striatum ”establishing and maintaining procedural habit structures”. (Wise 2004) Subsequently, the ventral striatum would correspond to the critic and the dorsal striatum to the actor of an RL algorithm (Sutton and Barto 2018). Dopamine would then correspond to the TD-error which is used to update both the actor and the critic. As dopamine neuron axons target both the ventral and dorsal striatum, and dopamine appears to be critical for synaptic plasticity, the similarities are evident. Furthermore, the TD-error and dopamine levels are both encoding the RPE: they are high whenever an unexpected reward is received and they are low (or negative in case of the TD-error) when an expected reward does not occur. (Sutton and Barto 2018) These similarities could be beneficial for RL as well as for Neuroscience as advances in either field could lead to new insights that are beneficial to the other.
Appendix G Timing comparison
In order to compare wall-clock time, we did a quick performance test for RTRRL with CT-RNN and PPO with LSTM. In both cases we trained for 5000 steps, with 32 units and a batch size of 1, on a machine with a single GeForce RTX 2070 GPU. For both cases, we repeated the test 3 times and calculated the average time per step of the algorithm. Our results were 7,58 ms / step for PPO-LSTM and 7,49 ms / step for RTRRL-RFLO. Please note that these results may not carry over to larger model or batch sizes.